OpenStudy (anonymous):

solve for x 83 = 33(to the x power)

OpenStudy (fibonaccichick666):

so how do you undo a power to start?

OpenStudy (fibonaccichick666):

OR what is a way to bring an exponent down?

OpenStudy (aum):

if \(\large a = b^x\), then \(\large \log(a) = \log(b^x) = x * \log(b)\) \(\large x = \Large \frac{\log(a)}{\log(b)}\)

OpenStudy (anonymous):

1.26?